# PoliCon: Evaluating LLMs on Achieving Diverse Political Consensus Objectives

### v2

@@ 4.1 Can Our Evaluator Simulate the Voting Results Well?
As mentioned in subsection 3.1, to construct our open-ended evaluation for assessing whether the tested LLM can reach the objectives of each political consensus, we first need a sub-evaluation module to score the preferences of each stakeholder. To this end, we introduce a GPT-4o-mini-backboned evaluator and request it to output an integer scalar between 0 and 9 to simulate the percentage of MEPs from each party who voted in favor.
@@
For evaluating its reliability, we randomly sample 100 issues for every combination of party and topic in each parliamentary term, resulting in approximately 41,800 testing samples in total. We conduct a consistency validation experiment by comparing our evaluator’s scores with the real-world party voting results, where we achieve a high Pearson correlation coefficient of 0.83.
Additionally, referring to existing work [Zhou et al., 2023], we plot Figure 4 to further illustrate the distribution of computational errors for our evaluator. The error is calculated by subtracting the ground truth voting score from the simulated score of our evaluator. It can be observed that the majority of the simulation results (>72%) are centered within the standard deviation $\sigma$ ($\pm$1.90) around the real voting results, with more simulation results showing a slight overestimation than underestimation. Based on the above experiments, we answered the question raised in the caption of the subsection with our evaluator is sufficiently capable of simulating each party’s voting results for the current resolution. More detailed experimental results can be found in Appendix E.1.
4.2 Performance Analysis for Various LLMs
@@ 4.3 Performance Analysis for Different Task Settings and Issue Topics
Analysis for Different Issue Topics.
As shown in 8(a), we analyze the experimental results of five coarse-grained topics. These results suggest that the difficulty of different topics shows certain similarities across various parliamentary settings. Specifically, topics involving policies, such as Security and Civil Rights, tend to be more challenging than those related to industrial development. This may be because these topics tend to present more complex and conflicting positions, requiring the evaluated LLM to possess stronger reasoning capabilities. For the complete experimental results of each fine-grained topic, see Appendix E.2.
4.4 Bias Evaluation for the Tested LLMs
@@ 5 Conclusion
In this work, we introduced PoliCon, a novel benchmark constructed from 2,225 European Parliament deliberation records to evaluate LLMs’ ability to achieve diverse political consensus objectives across diverse real-world collective decision-making settings. Our framework incorporates key factors like political issues, goals, party stances, and power structures, with an evaluation system that first simulates real voting outcomes and then assesses whether relevant political consensus can be achieved based on the social choice theory. Our experiments highlight PoliCon’s promise as an effective platform for studying LLMs’ ability to promote political consensus. To our knowledge, PoliCon represents the first comprehensive benchmark for assessing diverse political consensus achieving capabilities in LLMs, offering both a valuable evaluation platform and new insights into how LLMs navigate complex governance scenarios.
@@
The PoliCon benchmark is constructed from openly available sources, including the European Parliament website, HowTheyVote, and the VoteWatch Europe dataset. Both the official website of the European Parliament and HowTheyVote allow the use of their data as long as the source is cited, while the VoteWatch Europe dataset follows the CC 4.0 license. Importantly, while PoliCon is built upon authentic deliberation records, it does not reflect or predict the actual positions of the European Parliament, but rather serves as a research framework. Given the potential societal risks of applying AI in governance, such as reinforcing systemic biases, creating ideological echo chambers, fostering over-reliance on automated decision-making systems, and amplifying politically sensitive or divisive content, we urge all users of this benchmark to be acutely aware of these risks and to proceed with a high degree of caution and ethical responsibility. To further mitigate risks of potential misuse, we will release PoliCon under a research-only license agreement. The authors bear no responsibility for misuse or politically motivated interpretations. A more detailed ethical statement is provided in Appendix G.
@@ A.1.1 Data Sources
The data collection process for the PoliCon begins with the VoteWatch Europe dataset [HIX et al., 2022], which contains structured voting records of the European Parliament (EP) spanning 18 years from 2004 to 2022. Since the data for the five years from 2004 to 2009 is incomplete, we have excluded it. The portion of the dataset we use includes: (1) Excel files with metadata for the seventh, eighth, and half of the ninth European Parliament terms (2014-2022), including vote identifiers, titles, issue topics, etc.; (2) Roll call voting records mapping MEPs to vote outcomes, including six categories: in favor, against, abstain, absent, not voted, not an MEP; (3) URLs of the original sources from the official website of the European Parliament regarding where to obtain the voting data.
@@ A.5 Data Filtering Analysis
Final Valid Records 1,050 (15.1%) 1,127 (11.0%) 558 (4.1%)
@@
Final Valid Records.
The final valid records are those that are available and pass all quality checks. They are calculated by subtracting duplicates and missing records from the initial records. For example, in the 7th term, the calculation is 6,963 (initial records) - 5,333 (duplicates) - 580 (missing records) = 1,050 valid records. Despite the initial growth of the records, the number of valid records in the 9th term (558) decreased by 11% compared to the 7th term (1,050), highlighting the decline in the usability of the data.
@@ A.6 Data Cleaning Details
Due to the redundancy of the raw data, such as the large number of useless remarks in the debate, after collecting the original data, we further used DeepSeek-R1 [Guo et al., 2025] and rule-based methods for data cleaning and post-processing operations. First, we used DeepSeek-R1 to reorganize the resolutions and remove redundant parts while retaining the original resolution format. We further summarized the background of the current parliamentary discussion topics based on issue, resolution, and debate information.
@@
Appendix B Task Details
B.1 Topic Contents
Economics.
Industry.
Budget.
Security.
Civil Rights.
@@ B.2 More Stances Sematic Representation Results
As shown in Figure 9, Figure 10, and Figure 11, we present the sampled stances and all stances of all political parties during the 7th, 8th, and 9th terms of the parliament. From these three figures, it can be observed that the distribution results after sampling 200 data points for each party closely resemble those of the entire dataset, providing a strong reference value. Additionally, we can see that the distribution of party stances in the 7th and 8th terms of the European Parliament is more diverse compared to the 9th term. This may be due to factors such as Brexit [Besselink et al., 2019] and the rise of right-wing forces [Mudde, 2019, Servent, 2019, Abou-Chadi and Wagner, 2021], which highlights that our data analysis aligns with actual political trends.
@@ C.2 Task Prompts
The system prompt displayed above demonstrates good performance and has been widely used in previous work [Zhang et al., 2025].
@@ C.3 Evaluation Prompts
The experiments in subsection 4.1 and Appendix E.1 demonstrate a strong consistency between our evaluation method and the real voting results.
@@ Appendix D Quality of the Cleaned Data
Appendix E More Experimental Results
@@
E.1 Detailed Simulated Evaluation Consistency Results
@@
The error is calculated by subtracting the ground truth voting score from the simulated score of our sub-evaluation module. It can be observed that for each term of the parliament, most of our simulation results fall within the ground truth’s $\sigma$ range: 70% for the 7th, 70% for the 8th, and 77% for the 9th. From this more detailed error analysis, we can also see that in almost every parliament, our evaluator tends to overestimate slightly rather than underestimate, particularly evident in the 8th parliament. This may be related to factors such as the sycophancy of LLMs [Sharma et al., 2023, Kran et al., 2025].
E.2 Detailed Fine-grained Topics Results

### v3

@@ 4.1 Can Our Evaluator Simulate the Voting Results Well?
As mentioned in subsection 3.1, to construct our open-ended evaluation for assessing whether the tested LLM can reach the objectives of each political consensus, we first need a sub-evaluation module to score the preferences of each stakeholder. To this end, we introduce a GPT-4o-mini-backboned evaluator and request it to output an integer scalar between 0 and 9 to simulate the percentage of MEPs from each party who voted in favor. For evaluating its reliability, we will introduce it from two perspectives in the following: the consistency with real-world ground truth voting results and the consistency with human annotators.
@@
Consistency with real-world voting results.
For this experiment, we randomly sample 100 issues for every combination of party and topic in each parliamentary term, resulting in approximately 41,800 testing samples in total. We conduct a consistency validation experiment by comparing our evaluator’s scores with the real-world party voting results, where we achieve a high Pearson correlation coefficient of 0.83.
Additionally, referring to existing work (Zhou et al., 2023), we plot Figure 4 to further illustrate the distribution of computational errors for our evaluator. The error is calculated by subtracting the ground truth voting score from the simulated score of our evaluator. The detailed calculation process is shown in the Appendix E.1. It can be observed that the majority of the simulation results (>72%) are centered within the standard deviation $\sigma$ ($\pm$1.90) around the real voting results, with more simulation results showing a slight overestimation than underestimation. Based on the above experiments, we answered the question raised in the caption of the subsection with our evaluator is sufficiently capable of simulating each party’s voting results for the current resolution. More detailed experimental results can be found in Appendix F.1.
Consistency with human annotators.
In the above experiments, we demonstrate that our evaluator maintains a high level of agreement with real-world ground-truth voting results. However, to further verify that our evaluator can generalize to assessing consensus decisions generated by LLMs, we conduct an additional human study to examine whether it can maintain consistency with human annotators.
Because our task involves both AI and political science, we recruit 20 human annotators. Half of them have AI-related backgrounds, and the other half are law students, allowing the team to cover a broader range of relevant domain knowledge.
We further design and create a refined and user-friendly user interface (UI) to make the annotation process easier for human annotators. The page includes the following design elements: first, it presents the annotators with the issue title, relevant background information, party stances, and the corresponding generated resolution. Then, the annotators are asked to assign scores for the consistency between the resolution and the stances, as well as for the feasibility of the resolution. In the scoring section, we provide detailed criteria that explain the standards for each score range. Meanwhile, to accommodate annotators with different native languages, we also implemented a multilingual switching mode. The detailed UI design can be found in Appendix E.2.
@@
Table 2: Consistency of our evaluator with ground truth and human evaluations.
Ground truth Human eval.
Mean error 1.36 1.61
$\sigma$ 1.90 1.92
Within $\pm\sigma$ 72% 72%
We randomly sample 100 consensus resolutions generated by LLMs for evaluation. To ensure annotation robustness, each data item is assigned to three different annotators, yielding a total of 300 annotation results. Consistent with the evaluator’s scoring method, we use the expected values of consistency and feasibility as the final evaluation scores. As shown in Table 2, we present the consistency results of our evaluator compared with the ground truth and human evaluations. We report the mean error, the standard deviation $\sigma$, and the proportion of samples whose errors are within $\pm\sigma$.
It can be seen that the mean error between our evaluator and human annotations is only 1.61, and more than 72% of the data errors are within $\pm 1.92$, which demonstrates that our evaluator is sufficient to generalize in evaluating LLM-generated consensus resolutions. Another interesting phenomenon is that the consistency between our evaluator and human annotators is highly aligned with the consistency between our evaluator and the real-world ground-truth voting results, indicating that our consistency-computation method based on real voting data is also effective.
4.2 Performance Analysis for Various LLMs
@@
We further report results for two rule-based baselines, Random and Greedy. Random attains the lowest scores on all tasks, indicating that current LLMs have a nontrivial advantage over naive random choice in political consensus finding. Greedy generally outperforms Random but remains below LLM-generated consensus texts, except in the two-party passing resolution settings. In the SM and 2/3M settings, it surpasses Qwen2.5-32B, Llama-3.3-70B, and Qwen2.5-72B and ranks behind GPT-4o, and in the VP setting, it is comparable to Llama-3.3-70B. These results show that, under these conditions, most open-source models still retain substantial room for improvement.
@@ 4.3 Performance Analysis for Different Task Settings and Issue Topics
Analysis for Different Issue Topics.
As shown in 8(a), we analyze the experimental results of five coarse-grained topics. These results suggest that the difficulty of different topics shows certain similarities across various parliamentary settings. Specifically, topics involving policies, such as Security and Civil Rights, tend to be more challenging than those related to industrial development. This may be because these topics tend to present more complex and conflicting positions, requiring the evaluated LLM to possess stronger reasoning capabilities. For the complete experimental results of each fine-grained topic, see Appendix F.3.
4.4 Bias Evaluation for the Tested LLMs
@@ 5 Conclusion
In this work, we introduced PoliCon, a novel benchmark constructed from 2,225 European Parliament deliberation records to evaluate LLMs’ ability to achieve diverse political consensus objectives across diverse real-world collective decision-making settings. Our framework incorporates key factors like political issues, goals, party stances, and power structures, with an evaluation system that first simulates real voting outcomes and then assesses whether relevant political consensus can be achieved based on the social choice theory. Our experiments highlight PoliCon’s promise as an effective platform for studying LLMs’ ability to promote political consensus. To our knowledge, PoliCon represents the first comprehensive benchmark for assessing diverse political consensus achieving capabilities in LLMs, offering both a valuable evaluation platform and new insights into how LLMs navigate complex governance scenarios. We believe our work points to one of the important directions for the intersection research between AI and political science.
@@
The PoliCon benchmark is constructed from openly available sources, including the European Parliament website, HowTheyVote, and the VoteWatch Europe dataset. Both the official website of the European Parliament and HowTheyVote allow the use of their data as long as the source is cited, while the VoteWatch Europe dataset follows the CC 4.0 license. Importantly, while PoliCon is built upon authentic deliberation records, it does not reflect or predict the actual positions of the European Parliament, but rather serves as a research framework. Given the potential societal risks of applying AI in governance, such as reinforcing systemic biases, creating ideological echo chambers, fostering over-reliance on automated decision-making systems, and amplifying politically sensitive or divisive content, we urge all users of this benchmark to be acutely aware of these risks and to proceed with a high degree of caution and ethical responsibility. To further mitigate risks of potential misuse, we will release PoliCon under a research-only license agreement. The authors bear no responsibility for misuse or politically motivated interpretations. A more detailed ethical statement is provided in Appendix H.
@@
Acknowledgments
The work of PoliCon was sponsored by the National Natural Science Foundation of China (62376013, 623B2003, 624B100026). This work was also supported by the Natural Science Foundation of Beijing (QY24041). Any opinions, findings, or conclusions expressed in this work do not necessarily reflect the views of the funding agencies.
@@ A.1.1 Data Sources
The data collection process for the PoliCon begins with the VoteWatch Europe dataset (HIX et al., 2022), which contains structured voting records of the European Parliament (EP) spanning 18 years from 2004 to 2022. Since the data for the five years from 2004 to 2009 is incomplete, we have excluded it. The portion of the dataset we use includes: (1) Excel files with metadata for the seventh, eighth, and half of the ninth European Parliament terms (2014-2022), including vote identifiers, titles, issue topics, etc.; (2) Roll call voting records mapping MEPs to vote outcomes, including six categories: in favor, against, abstain, absent, not voted, not an MEP; (3) URLs of the original sources from the official website of the European Parliament regarding where to obtain the voting data.
@@ A.5 Data Filtering Analysis
Final Raw Valid Records 1,050 (15.1%) 1,127 (11.0%) 558 (4.1%)
@@
Final Raw Valid Records.
The final raw valid records are those that are available and pass all quality checks before the data cleaning process. They are calculated by subtracting duplicates and missing records from the initial records. For example, in the 7th term, the calculation is 6,963 (initial records) - 5,333 (duplicates) - 580 (missing records) = 1,050 valid records. Despite the initial growth of the records, the number of valid records in the 9th term (558) decreased by 11% compared to the 7th term (1,050), highlighting the decline in the usability of the data.
@@ A.6 Data Cleaning Details
Due to the redundancy of the raw data, such as the large number of useless remarks in the debate, after collecting the original data, we further used DeepSeek-R1 (Guo et al., 2025) and rule-based methods for data cleaning and post-processing operations. First, we used DeepSeek-R1 to reorganize the resolutions and remove redundant parts while retaining the original resolution format. We further summarized the background of the current parliamentary discussion topics based on issue, resolution, and debate information.
@@
After this data cleaning process, we once again inspected the data and carried out two filtering operations: (1) filtering out records with missing resolution or stances; and (2) classifying the final records by topic and removing topics with too few records. In the following paragraph, I will introduce these two steps separately:
Filtering records with Missing Resolution or Stances.
For records where the resolution is an empty string, the handling is straightforward by directly removing. The situation is slightly more complex for missing stances, because sometimes the stances string is not empty but is None or contains fewer than five words. Therefore, we adopt a rule-based method to filter original data with missing stances, namely, checking whether it is None, whether its length is less than 50 characters, or whether it is an empty string. If all stances of a record are filtered out based on these rules, the entire record is removed. In this process, removing empty resolutions filtered out 99 records, while removing empty stances filtered out 375 records. Thus, a total of 99 + 375 = 474 records were filtered out in this step.
Filtering Topics with Too Few records.
In this process, we first categorize all records according to their topics. Since some topics contain very few records, for example, the “juridical affairs” topic includes only 13 records, treating them as independent tasks would provide insufficient data. Therefore, we remove topics with 25 or fewer records. In this step, two topics are removed: “juridical affairs” (only 13 records) and “petitions” (only 23 records), resulting in a total of 36 filtered records.
Therefore, in these final two filtering steps, we filtered a total of $474+36=510$ records. After these two steps, the total number of records decreased from the original 2,735 to the final 2,225 complete ones.
@@
Appendix B Task Details
B.1 Topic Contents
Economics.
Industry.
Budget.
Security.
Civil Rights.
@@ B.2 More Stances Sematic Representation Results
As shown in Figure 9, Figure 10, and Figure 11, we present the sampled stances and all stances of all political parties during the 7th, 8th, and 9th terms of the parliament. From these three figures, it can be observed that the distribution results after sampling 200 data points for each party closely resemble those of the entire dataset, providing a strong reference value. Additionally, we can see that the distribution of party stances in the 7th and 8th terms of the European Parliament is more diverse compared to the 9th term. This may be due to factors such as Brexit (Besselink et al., 2019) and the rise of right-wing forces (Mudde, 2019; Servent, 2019; Abou-Chadi and Wagner, 2021), which highlights that our data analysis aligns with actual political trends.
@@ C.2 Task Prompts
The system prompt displayed above demonstrates good performance and has been widely used in previous work (Zhang et al., 2025).
@@ C.3 Evaluation Prompts
In our implementation, we use the average of the two metrics as the final score. The experiments in subsection 4.1 and Appendix F.1 demonstrate a strong consistency between our evaluation method and the real voting results.
@@ Appendix D Quality of the Cleaned Data
To further verify the impact of DeepSeek-R1 on our data-cleaning pipeline, we conduct a human evaluation experiment to measure the subjective quality of the data cleaned by DeepSeek-R1. We recruit ten human annotators and design an intuitive and easy-to-use interface for the experiment: the interface displays the original and cleaned resolutions and stances, and annotators are asked to evaluate the quality of the LLM-based data cleaning for each sample using a scalar score from 0 to 5. We randomly sample 100 instances from the dataset and assign 10 instances to each annotator as their labeling task. The results show that human annotators rate more than 95.8% of the cleaned resolution samples and more than 75% of the cleaned stances samples at 3 points or above. This indicates that DeepSeek-R1 is highly effective in improving data-cleaning quality and achieves an acceptable level of subjective quality. Details of the interface design appear in Appendix E.2.
Appendix E Experimental Details
In this section, we will introduce more details about how we implement our experiments.
E.1 Detailed Simulated Evaluation Consistency Calculation Process
We primarily used gaussian_kde from scipy.stats for error computation. Specifically, the complete computation process can be divided into the following three steps: (1) compute the errors; (2) compute the standard deviation of the errors; (3) apply Gaussian kernel density estimation (Gaussian KDE) to obtain a smooth estimate of the error distribution, which is then used for visualization. The detailed explanation of the formulas and procedures for each step is as follows:
Error Definition and Computation.
We define the error as the difference between the predicted score of our judge model and the real-world ground truth. The strict definition is as follows: suppose there are $n$ observed samples, and denote the set of sample indices as $\mathcal{I}={1,2,\dots,n}$. For each $i\in\mathcal{I}$, let $p_{i}\in[0,9]\cap\mathbb{N}$ be the predicted score, and $g_{i}\in[0,9]\cap\mathbb{N}$ be the corresponding ground-truth score, where $\mathbb{N}$ is the set of natural numbers. Then the prediction error of the $i$-th sample is defined as: $e_{i}=p_{i}-g_{i}$.
Standard Deviation of the Error.
We first compute the estimation of the errors $\bar{e}=\frac{1}{n}\sum_{i=1}^{n}e_{i}$, and then compute its standard deviation $\sigma=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(e_{i}-\bar{e})^{2}}$ with the estimation.
Using Gaussian KDE to Obtain and Visualize the Probability Density Function.
To obtain a smooth estimate of the error distribution, we apply Gaussian KDE to the empirical sample $\{e_{i}\}_{i=1}^{n}$, which is a method that treats each sample point as the center of a Gaussian distribution and forms a mixture of Gaussians to estimate the probability density function of a random variable. It is typically defined by the following formula:
$\hat{f}_{h}(x)=\frac{1}{nh\sqrt{2\pi}}\sum_{i=1}^{n}\exp\!\left(-\frac{1}{2}\left(\frac{x-e_{i}}{h}\right)^{2}\right),$ (2)
where $h>0$ is the bandwidth (smoothing) parameter. In this paper, we implement this function using gaussian_kde from scipy.stats, where $h$ is automatically determined by the Scott method (Scott, 2015) within gaussian_kde. Finally, we set $x\in[-9,9]$ with interval = $0.001$ for sampling and computation, thereby obtaining the error probability density curve shown in Figure 4.
E.2 Details for Our Human Study
Our human study can be divided into two main parts. The first part investigates whether the evaluations produced by our evaluator are consistent with the ratings given by human annotators. The second part examines the quality of the data cleaned by DeepSeek-R1. In the following sections, we provide a detailed description of our annotation pipeline and the user interface used in the study.
Does Our Evaluator Align with Human Annotators?
To verify whether our proposed evaluator can reliably score the consensus resolutions generated by LLMs, we design a concise and intuitive user interface that allows annotators to manually assess the degree to which model-generated resolution texts satisfy the party stances. As shown in Figure 12, the interface includes the following key components:
  • Problem Background and Task Information Display. The top of the interface shows the title of the current topic, relevant background information, and the participating political parties along with a summary of their positions, ensuring that annotators can make quality judgments based on a full understanding of the context.
  • Model-Generated Resolution Display. Annotators can view the full resolution text generated by the LLM and assess its consistency with the positions of each party, as well as the feasibility and reasonableness of the resolution content.
  • Scoring Area. To improve the reproducibility of scoring, we provide clear explanations of the score ranges (e.g., 0 to 9 corresponding to different levels of consistency) and require annotators to give separate scores for two dimensions: alignment, i.e., the extent to which the resolution accurately reflects and integrates the positions of various parties; and feasibility, i.e., whether the resolution is realistically implementable and logically coherent.
  • Multilingual Support. Considering that annotators may come from different native language backgrounds, we added a multilingual mode switch to the interface to improve usability and robustness of the annotation process.
Since our task involves the intersection of AI and political science, we recruited 20 human annotators, half of whom have an AI-related background and have obtained or are pursuing a PhD; the other half are law students, enabling coverage of a broader range of relevant domain knowledge.
We randomly selected 100 consensus resolutions generated by LLMs, and to ensure robustness of the evaluation, each sample was repeated three times and assigned to three different annotators. Using this UI, we obtained a total of 300 human annotations. Statistical results show that the average error between our evaluator and human annotators is only 1.61, with over 72% of sample errors falling within the $\pm 1.92$ range, indicating that the evaluator maintains good consistency with human judgments.
Quality of DeepSeek-R1 Data Cleaning.
To evaluate whether the structured content generated by DeepSeek-R1 during the data cleaning stage is sufficiently reliable, we further conducted a human quality assessment experiment. As shown in Figure 13, we designed a dedicated user interface for rating the quality of the cleaning results, which mainly consists of the following components:
  • Original and Cleaned Text Comparison Area. The interface clearly displays three types of information side by side: the original parliamentary resolution, the original party stances, and the cleaned data version generated by DeepSeek-R1. This side-by-side layout allows annotators to quickly judge whether the cleaning results are faithful, accurate, and free of missing important information.
  • Single-Dimension Quality Scoring Area. Annotators are required to assign an overall score from 0 to 5 for the cleaning quality of each sample, with explanations provided for the scores (e.g., 0 for severe errors, 5 for fully accurate).
Naturally, this UI also supports multilingual functionality, enabling displays in the native languages of different annotators. We recruited ten annotators and randomly selected 100 samples, assigning 10 samples to each annotator via the UI for independent annotation. The final experimental results indicate that human annotators rated over 95.8% of the resolution-cleaned samples and over 75% of the stance-cleaned samples with a score of 3 or higher. These findings demonstrate that DeepSeek-R1 is highly effective in ensuring the quality of cleaned data, achieving an acceptable, and in many cases, a favorable level in subjective quality assessments.
Figure 12: The UI for the human evaluation for our evaluator.
Figure 13: The UI for the data cleaning quality evaluation.
Appendix F More Experimental Results
@@
F.1 Detailed Simulated Evaluation Consistency Results
@@
The error is calculated by subtracting the ground truth voting score from the simulated score of our sub-evaluation module. It can be observed that for each term of the parliament, most of our simulation results fall within the ground truth’s $\sigma$ range: 70% for the 7th, 70% for the 8th, and 77% for the 9th. From this more detailed error analysis, we can also see that in almost every parliament, our evaluator tends to overestimate slightly rather than underestimate, particularly evident in the 8th parliament. This may be related to factors such as the sycophancy of LLMs (Sharma et al., 2023; Kran et al., 2025).
F.2 Impact of Party Seats on Veto Power
To examine how the number of seats held by each party affects the robustness of the task of veto power, we design two experiments. The first tests whether an increasing sample number in each party’s seats strengthens or weakens the robustness of veto power; the second tests how strongly different party seat allocations influence the robustness of the final outcome.
For the first experiment, since we randomly draw one seat allocation for each data point, computing each model’s overall expected outcome over a large set of data with varying seat distributions is clearly a robust way to test the model’s veto power. Therefore, if in an experiment where we sample multiple different numbers of seats for each issue and compute the corresponding expected outcomes, the resulting experimental findings are consistent with the performance of the models presented in Table 1, this would indicate that increasing the number of party seats can enhance the robustness of the veto power outcomes.
In one of our task scenarios, namely the development subtopic with 4 parties, we randomly generated 30 different party seat allocations for each data point and conducted experimental analysis on the veto power objective.
Table 10: The VP score with 30 samples for each issue on the ’development’ topic.
Model VP Score
Qwen2.5-32B $0.53\pm 0.04$
Llama-3.3-70B $0.52\pm 0.04$
Qwen2.5-72B $0.53\pm 0.03$
GPT-4o $0.56\pm 0.04$
Deepseek-V3.1 $0.57\pm 0.04$
Gemini-2.5 $0.59\pm 0.03$
As shown in Table 10, we present the mean and variance for different seat allocation outcomes. By examining the experimental results in the table, we find that they are entirely consistent with those reported in Table 1: Qwen2.5-32B, Llama-3.3-70B, and Qwen2.5-72B exhibit similar performance, which is lower than that of the three commercial models. For these three commercial models, this capability, from strongest to weakest, is: Gemini-2.5, Deepseek-V3.1, and GPT-4o. This finding corroborates our first experimental conclusion, namely that increasing the number of seats sample number held by a party will enhance, rather than diminish, the robustness of the veto power task.
For the second experiment, we can conduct the analysis by examining the variances in the table above. It can be seen that all the variances are kept within a controllable range. This result demonstrates that although different choices for the number of party seats do indeed affect the outcomes of the veto power task to some extent, the impact is not significant and remains within a relatively controllable range.
Combining the two experiments above, we arrive at the following two conclusions: (1) sampling different numbers of seats for each party can significantly enhance the robustness of the experimental results on veto power; and (2) the choice of specific party seat allocations has only a minor impact on the final results.
F.3 Detailed Fine-grained Topics Results